### v1

# Linear cost mutual information estimation and independence test of similar performance as HSIC

Abstract
Evaluation of statistical dependencies between two data samples is a basic problem of data science/machine learning, and HSIC (Hilbert-Schmidt Information Criterion) [1] is considered the state-of-art method. However, for size $n$ data sample it requires multiplication of $n\times n$ matrices, what currently needs $\sim O(n^{2.37})$ computational complexity [2], making it impractical for large data samples. We discuss HCR (Hierarchical Correlation Reconstruction) as its linear cost practical alternative of even higher dependence sensitivity in tests, and additionally providing actual joint distribution model by description of dependencies through features being mixed moments, starting with correlation and homoscedasticity, also allowing to approximate mutual information as just sum of squares of such nontrivial mixed moments between two data samples. Such single dependence describing feature is calculated in $O(n)$ linear time. Their number to test varies with dimension $d$ - requiring $O(d^{2})$ for pairwise dependencies, $O(d^{3})$ if wanting to also consider more subtle triplewise, and so on.
Keywords: independence test, joint distribution, entropy, mutual information, HSIC, HCR, normality test
I Introduction
Statistical dependencies start with correlation coefficients, defining dependence between expected values of two random variable: their first moments. There are also higher moments, like variance, skewness, kurtosis - which also contribute to statistical dependencies, like homoscedasticity evaluating co-occurrence of high variance e.g. in financial time series.
Figure 1: Proposed HCR independence test and comparison with state-of-art HSIC having much higher complexity, impractical for large samples. We generate shown $n=1000$ size random dataset from Gaussian mixtures independently for both coordinates, and slightly rotate it to introduce dependence. In HCR we first separately normalize coordinates to nearly uniform distribution in $[0,1]$ with CDF/EDF, then model joint density as a chosen degree polynomial, here $m=4$. Using orthonormal basis of Legendre polynomials, coefficients are approximately: expected value, variance, skewness, kurtosis, and their dependencies. Sum of squares of shown $m^{2}$ nontrivial coefficients approximate mutual information, and allows for independence test: by first normalizing multiplying by $\sqrt{n}$, and then testing if they agree with $N(0,1)$ set of variables, where the presented test of extreme coefficients has turned out the most sensitive to dependencies. There is shown probability of exceeding these values with min/max from $m^{2}=16$ independent $N(0,1)$ variables, also both by taking their product using shown Mathematica code, properly indicating dependence with much higher certainty than HSIC, additionally providing its description with mixed moments and model of joint density.
Hierarchical Correlation Reconstruction (HCR, [3, 4]) automatically decomposes dependencies into such mixed moments of two or more variables. Specifically, like in copula theory [5], we start with normalization of all variables to nearly uniform distribution in $[0,1]$ ($\sim$ quantiles) by $\hat{x}=\textrm{CDF}(x)$ or EDF (cumulative/empirical distribution function) separately for each coordinate. Then we model joint density of such normalized variables as a linear combination in product basis:
$\rho(\textbf{x})\approx\sum_{\textbf{j}\in B}a_{\textbf{j}}\,f_{\textbf{j}}(\textbf{x})=\sum_{(j_{1},\ldots,j_{d})\in B}a_{\textbf{j}}\,f_{j_{1}}(x_{1})\cdot\ldots\cdot f_{j_{d}}(x_{d})$ (1)
for bold $\textbf{x}=(x_{1},\ldots,x_{d})$ denoting dimension $d$ vectors, and $B$ chosen basis of mixed moments. For any continuous joint distribution we can approximate it with polynomials as close as we want, the $a_{\textbf{j}}$ coefficients become moments. Using orthonormal basis of polynomials $\int_{0}^{1}f_{k}(x)f_{l}(x)dx=\delta_{kl}$, coefficients can be MSE estimated as $a_{\textbf{j}}=\frac{1}{n}\sum_{\textbf{x}}f_{\textbf{j}}(\textbf{x})$ mean over normalized sample, in $O(n)$ time for $|X|=n$ data sample.
Thanks to orthonormality, we can approximate mutual information as sum of squares of such nontrivial moments between two variables [4]: $I(X,Y)\approx\sum_{\textbf{j},\textbf{k}\neq\textbf{0}}(a_{\textbf{j},\textbf{k}})^{2}$. As independent variables have zero mutual information, in this article we will adapt it for independence test, asking for probability if all tested nontrivial moments between two variables could be zero.
Alternative way to test independence is verifying if we can factorize joint distribution: $P_{X,Y}=P_{X}P_{Y}$?. It is used e.g. in HSIC independence test [1], considered as the state-of-art. However, it requires matrix multiplication for $n\times n$ matrices like $K_{ij}=\exp(-\|\textbf{x}^{i}-\textbf{x}^{j}\|_{2}/2\sigma^{2})$, and the lowest complexity known algorithm for matrix multiplication has $\sim O(n^{2.37})$ complexity [2], making this test impractical for large data samples.
Therefore, we discuss replacement of $\sim O(n^{2.37})$ HSIC independence test, with practical also for large samples: $O(n|B|)$ cost test based on HCR, for $B$ being a chosen set of features as mixed moments to test, e.g. growing $|B|\sim O(d^{2})$ with square of dimension if including only pairwise dependencies, or $O(d^{2})$ if adding triple-wise and so on.
The original motivation of this article ([4]) was application for information bottleneck training of neural networks [6], planned as future work. While originally it required mutual information evaluating the number of bits shared between contents of neural network layers, in practical realizations it was replaced with HSIC ([7, 8]), emphasizing it is something different, just practical to evaluate. In contrast, for HCR we indeed derive practical formula for approximation of mutual information, as just sum of squares of nontrivial mixed moments between two data samples, also using global basis which is often better for generalization as we can see in Fig. 2, hopefully leading to improvements of information bottleneck training.
This is initial version of article, with main purpose to introduce the method. We plan to extend benchmarks comparing with HSIC library [^1] HSIC library used for benchmarks: http://pypi.org/project/PyRKHSstats/ and details in the next version.
Figure 2: Comparison of joint density estimation using global basis of polynomials/moments we use in HCR and local basis of Gaussians usually used in HSIC, requiring to choose radius $\epsilon$. Global basis allows to represent the density with a reasonable number of features (moments) e.g. $m^{2}$ here, allowing the discussed complexity reduction. Evaluating log-likelihood in cross-validation, local basis barely got above 0 obtained also for trivial $\rho=1$ assumption, while global basis was able to find features as moments, which generalize well from training to test set. Here with the highest log-likelihood for $m=4$ degree up to kurtosis, we focus on in this article. Intuitively, global features like moments are better at generalization, than local features just assuming that new points will be close to the old points.
II HCR independence test and evaluation
Assume we have $Z=\{(\textbf{x}^{i},\textbf{y}^{i})\}_{i=1}^{n}$ as size $n$ data samples as vectors of $d_{x}$, $d_{y}$ dimensions, for which we would like to test independence. To simplify presentation we focus on $d_{x}=d_{y}=d$ equal dimension case, but also discuss general situation.
In HCR it is convenient to start with normalization of marginal distributions to nearly uniform in $[0,1]$, then model their joint distribution as a linear combination in some orthonormal basis, usually polynomials - which should use zero coefficients for independent variables, we will build independence test from.
II-A Normalization to nearly uniform in $[0,1]$ with CDF/EDF
As in copula theory citecopula, we start with normalization of variables to nearly uniform distribution in $[0,1]$ by transforming with CDF/EDF (cumulative or empirical distribution function) to nearly quantiles, independently for each coordinate.
For CDF approach we need to choose some parametric family e.g. Gaussian, and estimate its parameters e.g. based on dataset, separately for each coordinate. Generally, especially if data does not suit any parametric family, we can use EDF normalization: order values in the size $n$ sample and assign (i1/2)/n(i-1/2)/n to $i$-th in the order. However, it requires sorting having $O(n\ln n)$ complexity - if linear is required, there should be used CDF.
Figure 3: Top left: the first $j=0,1,2,3,4$ from orthonormal polynomial basis $f_{j}(x)$ for $d=1$ dimension and product basis $f_{j}(x)f_{k}(y)$ for $d=2$ dimensions, which linear combination is used as model of (joint) density in HCR, e.g. $\rho(x,y)=\sum_{jk=1}^{m}a_{jk}f_{j}(x)f_{k}(y)$ for variables normalized to nearly uniform distribution in $[0,1]$. As $f_{0}=1$, top row $a_{i0}$ describes marginal distribution of $x$ and $a_{0j}$ left column of $y$. Then $i,j\geq 1$ describe their dependencies as mixed moments close to expected value, variance, skewness and kurtosis. Right: $d=3$ case adding triplewise dependencies.
Bottom left: standard view on (joint/conditional) entropy of $X,Y$ variables and their mutual information. Right: for HCR using $\ln(1+w)\approx w$ first Taylor term they can be approximated with sums of squares of coefficients, especially mutual information as $I(X;Y)\approx\sum_{i,j=1}^{m}\left(a_{ij}\right)^{2}$, also in higher dimension, which should be close to zero for independent variables.
For the discussed independence test/evaluation in benchmarks we use EDF: for $d$-dimensional data samples $X$ and $Y$, we start with transforming them by applying EDF normalization separately for each coordinate as $\textrm{EDF}^{x}_{i},\textrm{EDF}^{y}_{i}$:
$\hat{Z}=\{(\hat{\textbf{x}},\hat{\textbf{y}}):\hat{x}_{i}=\textrm{EDF}^{x}_{i}(x_{i}),\hat{y}_{i}=\textrm{EDF}^{y}_{i}(y_{i}),(\textbf{x},\textbf{y})\in Z\}$ (2)
As mutual information is invariant under reparametrization of the marginal variables [9], we can estimate it for such transformed variables instead: $I(U;V)\approx I(X;Y)$.
II-B Polynomial density model
Let us now model joint distribution on $(\textbf{x},\textbf{y})\in\hat{Z}$ normalized variables as a linear combination of usually polynomials $f$:
$\rho(\textbf{x},\textbf{y})\approx\sum_{\textbf{j}\in B_{x},\textbf{k}\in B_{y}}a_{\textbf{j},\textbf{k}}\,f_{\textbf{j}}(\textbf{x})f_{\textbf{j}}(\textbf{y})$ (3)
for chosen bases: $B_{x}$$B_{y}$ (can be equal), e.g. $\{0,\ldots,m\}^{d}$. Using orthonormal family, MSE coefficient estimation becomes:
$a_{\textbf{j},\textbf{k}}=\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))=\frac{1}{n}\sum_{(\textbf{x},\textbf{y})\in\hat{Z}}f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y})$ (4)
where we can use product basis $f_{\textbf{j}}(\textbf{x})=f_{j_{1}}(x_{1})\cdot\ldots\cdot f_{j_{d}}(x_{d})$ for orthonormal: $\int_{0}^{1}f_{k}(x)f_{l}(x)dx=\delta_{kl}$ e.g. polynomials (rescaled Legendre) we will use, getting interpretation close to standard cumulants:
$f_{0}=1\qquad\textrm{corresponds to normalization}$ (5)
$f_{1}(x)=\sqrt{3}(2x-1)\qquad\sim\textrm{expected value}$
$f_{2}(x)=\sqrt{5}\left(6x^{2}-6x+1\right)\qquad\sim\textrm{variance}$
$f_{3}(x)=\sqrt{7}\left(20x^{3}-30x^{2}+12x-1\right)\qquad\sim\textrm{skewness}$
$f_{4}(x)=3\left(70x^{4}-140x^{3}+90x^{2}-20x+1\right)\qquad\sim\textrm{kurtosis}$
Alternatively there could be used Fourier or DCT basis, especially for periodic variables.
The coefficients can be interpreted as mixed moments, as $f_{0}=1$ only between nonzero indexes, e.g. $a_{00i0}$ is $i$-th moment of 3rd out of 4 variables, $a_{012}$ describe dependence between expected value of 2nd variable, and variance of 3rd variable.
II-C Entropy, mutual information estimation
Using $\ln(1+t)\approx t$ approximation and orthonormality, we can approximate entropy in nits ($1/\ln(2)\approx 1.44$ bits) from the coefficients:
$H(\hat{X})=-\int_{[0,1]^{d}}\rho(\mathbf{x})\,\ln(\rho(\mathbf{x}))\,d\mathbf{u}\approx-\sum_{\mathbf{j}\in B_{x}^{+}}(a_{\mathbf{j}})^{2}$ (6)
for $B\{$$^{+}=B\backslash\{\textbf{0}\}$ basis without normalization $f_{\textbf{0}}=1$. Like visualized in Fig. 3, analogously e.g. for joint distribution:
[…]
allowing to approximate mutual information as just sum of squares of nontrivial mixed moments between variables:
$I(\hat{X},\hat{Y})=H(\hat{X})+H(\hat{Y})-H(\hat{X},\hat{Y})\approx\sum_{\mathbf{j}\in B_{x}^{+}}\ \sum_{\mathbf{k}\in B_{y}^{+}}\left(a_{(\mathbf{j},\mathbf{k})}\right)^{2}$ (8)
However, using these sum of squares formulas with (4) estimator as mean: $a_{\textbf{j},\textbf{k}}=\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))$, it would be artificially increased - let us now discuss it and correction.
II-C1 Subtracting the variance
Assume we get $n$ values $\{w^{i}\}_{i=1..n}$ from $N(\mu,\sigma)$ distribution, and we want to estimate $\mu^{2}$. A natural approach is calculating the mean $\bar{w}=\frac{1}{n}\sum_{i}w_{i}$ and taking $\bar{w}^{2}$, however, it would be artificially increased.
The Central Limit Theorem (CLT) says that for large sample:
$\textrm{CLT}:\qquad\qquad\frac{\bar{w}-\mu}{\sigma/\sqrt{n}}\sim N(0,1)$ (9)
Applying square to both sides, and taking expected values: mean square from $N(0,1)$ is 1, hence we get:
$n\,E[(\bar{w}-\mu)^{2}/\sigma^{2}]=1$
As $E[\bar{w}]=\mu$, expanding the square we finally get
$\mu^{2}=E[\bar{w}^{2}]-\sigma^{2}/n$ (10)
[…]
$I_{c}(\hat{X},\hat{Y})=\sum_{\mathbf{j}\in B_{x}^{+}}\ \sum_{\mathbf{k}\in B_{y}^{+}}\left(\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))^{2}-\frac{1}{n}\textrm{var}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))\right)$ (11)
we plan to compare in future with state-of-art methods like [9].
II-D HCR-based independence tests
[…]
To test it, let us first normalize them as in CLT:
$M=|B^{+}_{x}||B^{+}_{y}|\qquad\textrm{of}\qquad\hat{a}_{\textbf{j},\textbf{k}}=\frac{\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))}{\sqrt{\textrm{var}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))/n}}$ (12)
Assuming $H_{0}$ hypothesis, for pairwise dependencies: between one coordinate of both samples, due to independence and normalization above ’var’ variance has to be 1, allowing to omit the above estimation from sample. However, for higher order dependencies it does not have to be true due to dependencies inside samples - requiring to estimate variance from sample.
Accordingly to CLT, for $H_{0}$ independence hypothesis, all $\hat{a}_{\textbf{j},\textbf{k}}$ should be from approximately $N(0,1)$.
[…]
II-D1 Log-likelihood test
of $\textrm{mean}(\ln(\textrm{PDF}(\hat{a}_{\textbf{j},\textbf{k}})))$, which assuming $H_{0}$ and large $M$ is approximately from normal distribution. For $N(0,1)$ PDF we get $E[\ln(\textrm{PDF})]=-(1+\ln(2))/2\approx-1.4189$ and variance 1/2, leading to test:
$\textrm{Pr}(H_{0})=1-\textrm{erf}\left(\frac{|\textrm{mean}(\ln(\textrm{PDF}(\hat{a}_{\textbf{j},\textbf{k}})))+1.4189|}{2\sqrt{M}}\right)$ (13)
Figure 4: $\textrm{PDF}^{M}_{i}$ probability densities of sorted $M=1,\ldots,16$ variables from $N(0,1)$ normalized Gaussian distribution, which can be used for tests if estimated normalized coefficients are from $N(0,1)$ for $\mathcal{H}_{0}$ hypothesis of independence, especially for the extreme values indicating the main dependence direction.
II-D2 Testing extremes $\{\hat{a}_{\textbf{j},\textbf{k}}\}_{\textbf{j}\in B_{x}^{+},\textbf{k}\in B_{y}^{+}}$
[…]
For hypothesis testing, we need to find distribution for $M$ sorted values from $N(0,1)$. To find $\textrm{PDF}_{M,i}(x)$ density of $i$-th in order, there are i1i-1 smaller or equal, and M1iM-1-i greater then equal, combinatorially leading to $M{M-1\choose i-1}$ possibilities, with density, shown in Fig. 4:
$\textrm{PDF}^{M}_{i}(x)=M{M-1\choose i-1}\,\textrm{PDF}(x)\,\textrm{CDF}(x)^{i-1}(1-\textrm{CDF}(x))^{M-i}$ (14)
for $\textrm{PDF}\equiv\textrm{PDF}^{1}_{1}$, $\textrm{CDF}\equiv\textrm{CDF}^{1}_{1}$ of $N(0,1)$. By integrating it we can get its CDF for values from $N(0,1)$:
$\textrm{CDF}^{M}_{i}(x)=M{M-1\choose i-1}\,\textrm{Beta}(\textrm{CDF},i,1-i+M)$ (15)
using incomplete Euler Beta function. We can use $\textrm{CDF}_{M,i}$ to test probability of obtaining one or a few extreme values.
II-E Choice of basis $B$
For $d_{x}=d_{x}=d=1$ dimensional samples, a natural choice of basis is just $B^{+}=\{1,2,\ldots,m\}$ for some $m$ number of moments, e.g. for $m=1$ we have only $a_{1,1}$ corresponding to just testing correlation. For $m=2$ we would additionally include homoscedasticity as $a_{2,2}$, but also $a_{1,2}$ and $a_{2,1}$ dependencies between expected value and variance, and so on, generally using $m^{2}$ features describing dependencies between data samples.
Examples in this article use $m=4$ up to kurtosis with $m^{2}=16$ features, which through cross-validation usually turn out optimal e.g. for financial data as in Fig. 2, can be optimized for some specific data this way, or e.g. trying to recognize small rotations for like in Fig. 1.
For larger $d\geq 1$ dimensions it usually should be sufficient to focus on $(md)^{2}$ pairwise dependencies for bases with single nonzero index:
$\textrm{pairwise:}\quad B=\left\{\textbf{j}\in\{0,\ldots,m\}^{d}:\sum_{i}\textrm{sign}(j_{i})=1\right\}$ (16)
for fixed degree $m$, e.g. $m=1$ would correspond to testing only $d^{2}$ correlations between all pairs of coordinates.
While we have discussed $d_{x}=d_{y}=d$, the two samples can be of different dimension, where we still can calculate mixed moments between their coordinates, estimate mutual information with (8) formula using different bases. Also, while for simplicity we have discussed fixed $m$, it might be worth to vary it between coordinates, maybe also change the basis e.g. to Fourier for periodic coordinates like day of year.
Dependence between samples should be usually detectable through such pairwise dependencies. However, in theory there can be only more subtle higher order dependencies - HCR independence test would need to include to be able to detect, evaluate, describe. For example by adding to $B_{x}$ and/or $B_{y}$ basis $\sum_{i}\textrm{sign}(j_{i})=2$ indexes, e.g. $a_{11,1}$ would include triplewise dependence like in Fig. 3: that with change of both coordinates, there is change of expected value in the third one. The cost is larger number of features: $O(d^{2})$ for pairwise, $O(d^{3})$ for triplewise, and so on.
We can also perform some preprocessing of data to try to emphasize dependencies, hopefully making it more likely for low order e.g. pairwise test to find them, like PCA (principal component analysis) - rotation of coordinates to eigenvectors of correlation matrix, separately for both samples. Also e.g. using SVD (singular value decomposition) of $a_{\textbf{k},\textbf{j}}$ as $|B^{+}_{x}|\times|B_{y}^{+}|$ matrix, we can optimize both used bases. We can freely perform such optimization working on a family of data samples of similar behavior. However, performing such optimization separately on a single tested sample would also amplify its noise - it would need modification of independence test to include it, e.g. based on random matrix theory, or synthetic empirical simulations.
II-F HCR independence test used in benchmarks
The benchmarks for HCR have used basic pairwise dependencies for fixed degree $m=4$, testing minimal and maximal considered $\hat{a}$, presented as Algorithm 1 and in Fig. 1.
Algorithm 1 HCR basic pairwise independence test()
$n,d_{x},d_{y},m\in\mathbb{N}$   {sample dimensions, chosen degree}
$U=\textrm{EDF}(X)$    {coordinate-wise normalizations}
$V=\textrm{EDF}(Y)$    {separately sorting coordinates}
$\hat{a}_{j,q,k,r}=\sqrt{n}\ \textrm{mean}(f_{j}(u_{q})f_{k}(v_{r}))$  {normalized features}
    for $q=1,\ldots,d_{x},r=1,\ldots,d_{y},$ and $j,k=1,\ldots,m$
$M=d_{x}d_{y}m^{2}$       {number of features}
[…]
Accepting $H_{0}$ hypothesis if returned Pr(independent) is above some chosen significance level, or reject otherwise.
It uses $\textrm{var}=1$ in (12) as mentioned allowed for pairwise dependencies, for triplewise and higher $\hat{a}$ should be additionally divided by estimated variance for coefficients $\textbf{j},\textbf{k}$ having two or more nonzero indexes. For pairwise there is single nonzero index - we have denoted its position as $q$ and $r$.
Its modifications can vary $m$ and $f$ bases between coordinates, add triplewise and higher order dependencies, include basis optimizations.
It is also worth to consider different final tests if all $M$ normalized features are from $N(0,1)$ distribution - testing min and max is very sensitive to dependencies, but could lead to false positives, and ignores intermediate values. It can be extended from $2$ to $2k$ extreme values by sorting $\hat{a}$ to $\textrm{Pr}(\textrm{independent})\approx$
$2^{2k}\prod_{i=1}^{k}\textrm{CDF}_{i}^{M}\left(\textrm{sort}(\hat{a})_{i}\right)\,\left(1-\textrm{CDF}_{M+1-i}^{M}\left(\textrm{sort}(\hat{a})_{M+1-i}\right)\right)$ (17)
Where $2^{2k}$ is rescaling to get approximately 1 for independent. To include all values we could use e.g. Anderson-Darling, Shapiro-Wilk tests, $\chi^{2}$, permutation, or log-likelihood test.
II-G Toward kernel methods: global HCR vs local HSIC basis
As in [4], we can represent the features as $n\times|B|$ matrices:
$\bar{X}=\frac{1}{\sqrt{n}}\left(f_{\textbf{j}}(\textbf{x})\right)_{\textbf{x}\in\hat{X},\,\textbf{j}\in B_{x}^{+}}\qquad\bar{Y}=\frac{1}{\sqrt{n}}\left(f_{\textbf{k}}(\textbf{y})\right)_{\textbf{y}\in\hat{Y},\,\textbf{k}\in B_{y}^{+}}$ (18)
This way MSE estimation becomes $a_{\textbf{j},\textbf{k}}=\bar{X}^{T}\bar{Y}$ matrix, allowing to approximate mutual information (as trace is cyclic):
$I(X,Y)\approx\textrm{Tr}(\bar{X}^{T}\bar{Y}(\bar{X}^{T}\bar{Y})^{T})=\textrm{Tr}(K^{X}K^{Y})$ (19)
$\textrm{for }n\times n:\qquad K^{X}=\bar{X}\bar{X}^{T}\qquad K^{Y}=\bar{Y}\bar{Y}^{T}$
kernel matrices we will also use for HSIC in the next section, with included correction like (11) variance subtraction.
However, while in HCR we focus on global basis of e.g. polynomials or Fourier, for HSIC there is usually used local basis like Gaussians, leading to essentially different types of represented density models, like in Fig. 2, with different behavior:
  • global basis is usually better at generalization: provides higher log-likelihood in cross-validation, as describing general features, instead of just assuming that new points will be close to the old points for local basis,
  • global basis often allows to describe density with a reasonable number of features (moments), allowing more practical description e.g. fo,x’)r the discussed reduced complexity,
  • local basis requires to choose kernel width, while global to choose basis with often universal e.g. $m=4$ degree,
  • global orthonormal basis for normalized variables can work on deformation of $\rho=1$, allowing for practical approximations of entropy, mutual information using $\ln(1+t)\approx t$.
  • above HCR kernel matrices include all $B_{x}\times B_{y}$ features, while for HSIC Gaussian kernel only single $\sigma$ is used:
$K^{\textrm{HCR}}_{\textbf{x},\textbf{x}^{\prime}}=\frac{1}{n}\sum_{\mathbf{j}\in B_{x}^{+},\mathbf{k}\in B_{x}^{+}}f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{x}^{\prime})\quad\qquad K^{\textrm{HSIC}}_{\textbf{x},\textbf{x}^{\prime}}=e^{-\frac{\|\textbf{x}-\textbf{x}^{\prime}\|_{2}}{2\sigma^{2}}}$
III Hilbert–Schmidt Independence Criterion (HSIC)
III-A Statistical dependence measures
Let $(X,Y)$ be random variables with joint distribution $\mathcal{R}$ on $\mathcal{X}\times\mathcal{Y}$. A (population) dependence measure is a functional
$\Psi:\ \mathcal{P}(\mathcal{X}\times\mathcal{Y})\to[0,\infty).$
Following classical desiderata (e.g., Rényi), one may require (see [12]):
  • $\Psi$ is well-defined
  • $0\leq\Psi(\mathcal{R})\leq 1$
  • $\Psi(\mathcal{R})=0$ if and only if $X,Y$ are independent
  • $\Psi(\mathcal{R})=1$ if and only if $Y=g(X)$ for some deterministic bijective $g$
These are optional properties and do not have to hold for every measure; in particular, HSIC is nonnegative and equals zero iff $X$ and $Y$ are independent (with characteristic kernels), but it is not generally bounded by $1$ and does not attain $1$ for deterministic dependence.
A standard example of such statistical dependence measure in mutual information measuring the number of bits one variable says about the second, however, changing the upper bound from 1 to entropy for samples differing by bijection. In contrast to HCR, HSIC does not have such interpretation, however, is used as its practical replacement e.g. in information bottleneck training ([7, 8]).
In practice we work with a sample $Z=\{(\textbf{x}^{i},\textbf{y}^{i})\}_{i=1}^{n}$ and a sample statistic $\widehat{\Psi}:(\mathcal{X}\times\mathcal{Y})^{n}\to\mathbb{R}$ used as a test statistic.
III-B Definitions
III-B1 Reproducing Kernel Hilbert Spaces
Let $\mathcal{X}$ be set the data comes from and $\mathcal{H}$ a Hilbert space. Let $\varphi:\mathcal{X}\rightarrow\mathbb{R}$ be the feature map, mapping datapoints to points in $\mathcal{H}$.
We call $\kappa:\mathcal{X}\times\mathcal{X}\rightarrow\mathbb{R}$ a kernel if [13]
$\kappa(\textbf{x},\textbf{y})=\langle\varphi(\textbf{x}),\varphi(\textbf{y})\rangle_{\mathcal{H}}$
We call $\kappa$ a reproducing kernel and the Hilbert space $\mathcal{H}$ of functions $\mathcal{X}\rightarrow\mathbb{R}$ a reproducing kernel Hilbert Space if the following conditions are satisfied for every $x\in\mathcal{X}$ and every $f\in\mathcal{H}$:
  1. 1.
    $\kappa(\textbf{x},.)\in\mathcal{H}$
  2. 2.
    $f(\textbf{x})=\langle f,\kappa(\textbf{x},.)\rangle_{\mathcal{H}}$
The most popular kernel choice is the Gaussian function, so
[…]
for some fixed variance $\sigma$.
In this case $\mathcal{H}$ is the space of Gaussian mixtures and
[…]
III-B2 Kernel Mean Embedding
Let $\kappa$ be a reproducing kernel and $\mathcal{H}$ a RKHS. Let $\mathcal{P}:\mathcal{X}\rightarrow\mathbb{R}$ be a probability distribution on $\mathcal{X}$. Then, by the Riesz Representation Theorem, there exists an unique $\mu_{P}\in\mathcal{H}$ such that $\forall f\in\mathcal{H}$
$\langle f,\mu_{P}\rangle_{\mathcal{H}}=\mathbb{E}_{\textbf{x}\sim\mathcal{P}}(f(\textbf{x}))$
It can be written as [13]
$\mu_{P}=\mathbb{E}_{x\sim\mathcal{P}}(\kappa(x,.))$
III-B3 Cross-Covariance Operator
Let $X:\mathcal{X}\rightarrow\mathbb{R},Y:\mathcal{Y}\rightarrow\mathbb{R}$ be two random variables. Let $\mathcal{F},\mathcal{G}$ be their respective RKHS’s with associated kernel functions $k,l$. Let $\mathcal{R}$ be the joint distribution of $X,Y$ and $\mathcal{P}$, $\mathcal{Q}$ associated marginal distributions.
The cross-covariance operator $C_{\mathcal{R}}:\mathcal{F}\rightarrow\mathcal{G}$ is defined by satisfying the following condition for every $f\in\mathcal{F},g\in\mathcal{G}$:
$\langle f,C_{\mathcal{R}}(g)\rangle_{\mathcal{F}}=\mathbb{E}_{\textbf{x},\textbf{y}:\mathcal{R}}(f(\textbf{x})-\mathbb{E}_{\textbf{x}^{\prime}:\mathcal{P}}(f(\textbf{x}^{\prime})))(g(y)-\mathbb{E}_{\textbf{y}^{\prime}:\mathcal{Q}}(f(\textbf{y}^{\prime})))$
Which can be written as [13]
$C_{\mathcal{R}}=\mathbb{E}_{\textbf{x},\textbf{y}\sim\mathcal{R}}(k(\textbf{x},.)-\mu_{\mathcal{P}})\otimes(l(\textbf{y},.)-\mu_{\mathcal{Q}}))$
where $\otimes$ denotes a tensor product of these two Hilbert spaces.
III-B4 Hilbert-Schmidt Norm
Let $\mathcal{F},\mathcal{G}$ be Hilbert spaces and $f\in\mathcal{F},g\in\mathcal{G}$. Then the Hilbert-Schmidt norm of the tensor product $f\otimes g$ is defined as
[…]
III-C HSIC Formula
Value of the HSIC measure of the joint probability distribution $\mathcal{R}$ is the Hilbert-Schmidt norm of its covariance operator, so $C_{\mathcal{R}}||_{HS}$. This can be further transformed into [13]:
$\|C_{\mathcal{R}}\|_{HS}=\mathbb{E}_{\textbf{x},\textbf{y}\sim\mathcal{R},\textbf{x}^{\prime},\textbf{y}^{\prime}\sim\mathcal{R}(k(\textbf{x},\textbf{x}^{\prime})l(\textbf{y},\textbf{y}^{\prime}))+}$ (20)
$\mathbb{E}_{\textbf{x}\sim\mathcal{P},\textbf{x}^{\prime}\sim\mathcal{P}(k(\textbf{x},\textbf{x}^{\prime}))\mathbb{E}_{\textbf{y}\sim\mathcal{Q},y^{\prime}\sim\mathcal{Q}(l(\textbf{y},\textbf{y}^{\prime}))-2\mathbb{E}_{\textbf{x},\textbf{y}\sim\mathcal{R}}(\mathbb{E}_{\textbf{x}^{\prime}\sim\mathcal{P}}(k(\textbf{x},\textbf{x}^{\prime}))\mathbb{E}_{\textbf{y}^{\prime}\sim\mathcal{Q}}(l(\textbf{y},\textbf{y}^{\prime})))}}$
III-D HSIC Estimators
In practice stastistical dependence tests are performed on a set of $n$ datapoints instead of the distribution formula. Let
$Z=\{(\textbf{x}^{1},\textbf{y}^{1}),(\textbf{x}^{2},\textbf{y}^{2}),\ldots,(\textbf{x}^{n},\textbf{y}^{n})\}$
be the dataset.
Let $K,L\in\mathbb{R}^{n\times n}$ be kernel matrices such that
$K_{ij}=k(\textbf{x}^{i},\textbf{x}^{j})\qquad L_{ij}=l(\textbf{x}^{i},\textbf{x}^{j})$ (21)
The Gaussian function is typically chosen as $k,l$.
Let
$H=I_{n}-\frac{1}{n}J_{n}$
be the centering matrix where $J_{n}$ is an $n\times n$ matrix with all ones.
Then the V-statistic for HSIC is [13, 1, 12]:
$\textrm{HSIC}(Z)=\frac{1}{n^{2}}\textrm{Tr}(KHLH)$
III-E HSIC as a stastical test
In order to turn HSIC into the statistical dependence test, we state the null hypothesis $H_{0}$ as $\mathcal{R}=\mathcal{P}\mathcal{Q}$ and the alternative hypothesis $H_{1}$ as $\mathcal{R}\neq\mathcal{P}\mathcal{Q}$. Given the test statistic $\textrm{HSIC}(Z)$ and the desired test level $\alpha$, null hypothesis is rejected if $\textrm{HSIC}(Z)>\theta$ for some threshold $\theta$ such that the probability of type I error is $\alpha$.
The most popular method for deriving $\theta$ is approximating the distribution of $\textrm{HSIC}(Z)$ under $H_{0}$ with Gamma function [1]
$\textrm{HSIC}(Z)\approx\frac{x^{a-1}e^{-\frac{x}{b}}}{b^{a}\Gamma(a)}$
where
$a=\frac{\mathbb{E}(\textrm{HSIC}(Z))^{2}}{\textrm{var}(\textrm{HSIC}(Z))}$
$b=\frac{n\cdot\textrm{var}(\textrm{HSIC}(Z))}{\mathbb{E}(\textrm{HSIC}(Z))}$
Then $\theta$ is equal to its $1-\alpha$ quantile.
III-F HSIC vs. HCR
Both HSIC (Hilbert–Schmidt Independence Criterion) and HCR (Hierarchical Correlation Reconstruction) produce sample statistics that increase with dependence, but they arise from different geometries and offer complementary trade-offs.
Geometry and what is being measured
HSIC operates in an RKHS induced by user-chosen kernels $k$ and $\ell$ on $\mathcal{X}$ and $\mathcal{Y}$. With a characteristic choice (e.g., RBF), HSIC is zero if and only if $X$ and $Y$ are independent. Equivalently, HSIC is the squared MMD between the joint and the product of marginals with product kernel $s((x,y),(x^{\prime},y^{\prime}))=k(x,x^{\prime})\,\ell(y,y^{\prime})$. HCR works in a global, orthonormal polynomial basis on the copula domain after marginal normalization (CDF/EDF). Mixed moments $a_{\mathbf{j},\mathbf{k}}$ capture interpretable interactions (mean–mean, mean–variance, etc.). A practical MI proxy is the sum of squares of nontrivial mixed moments.
“Local” vs. “global” sensitivity
RBF kernels do not impose a KDE model per se; rather, the bandwidth $\sigma$ controls the effective locality of the RKHS features. For characteristic kernels, HSIC can capture global dependencies; poor bandwidth selection may reduce power to either very local or overly smooth effects. In HCR, the polynomial degree $m$ controls global complexity: higher $m$ adds higher-order mixed moments (e.g., skewness, kurtosis interactions). This often yields good out-of-sample generalization for global structure and offers straightforward interpretability.
Computational profile
A standard HSIC implementation forms $n\times n$ Gram matrices, incurring $O(n^{2})$ time and memory. This is often the practical bottleneck for large $n$. HCR computes each feature in one pass, $O(n)$, so testing a set $B$ mixed-moment features costs O(n|B|)O(n\,|B|) time and $O(|B|)$ memory, which scales well to large $n$ (with linear streaming updates).
Calibration and hyperparameters
HSIC thresholds can be obtained via permutation or gamma moment-matching for the biased V-statistic; the main hyperparameter is kernel choice and bandwidth (often tuned by CV or median heuristic). HCR requires choosing the basis (e.g., Legendre) and degree $m$, and possibly which coordinates/pairs to include. Calibration can use normalization of features to $N(0,1)$ under $H_{0}$ and multiple-testing control (e.g., via the distribution of $\max_{i}|\hat{z}_{i}|$ or permutations).
Interpretability and extensibility
HSIC yields a single scalar without a direct decomposition into human-readable effects. HCR is explicitly decomposable: each coefficient corresponds to a mixed moment, enabling diagnosis of which interactions drive dependence. Extending from pairwise to triplewise or higher-order interactions is straightforward by enlarging $B$ (with cubic and higher growth in $d$ if needed).
When to prefer which
For small/medium $n$ and when a powerful, kernel-agnostic test is sufficient, HSIC is a strong default. For very large $n$, streaming settings, or when interpretability and targeted feature testing matter, HCR is attractive due to linear cost and mixed-moment readouts. In practice, they are complementary: HSIC for broadly powered detection, HCR for scalable detection plus explanation of the dependence structure.
Figure 5: HSIC (Gamma) vs HCR. We start from $n=1000$ i.i.d. samples $(x_{1},x_{2},y_{1},y_{2})$ with $X=(x_{1},x_{2})$ independent of $Y=(y_{1},y_{2})$. A $50^{\circ}$ within–block rotation is applied to $(x_{1},x_{2})$ (which preserves independence), and then a fixed global rotation $R\in\mathbb{R}^{4\times 4}$ is applied repeatedly to the whole 4D vector; each application adds $\Delta\theta=0.25^{\circ}$ (the $x$-axis shows the cumulative angle). At each angle we test $H_{0}\!:\,X\perp Y$ at $\alpha=0.05$ and plot $p$-values on a log scale (red dashed line marks $\alpha$). HSIC (Gamma) uses Gaussian RBF kernels with bandwidths set by the median heuristic on $X$ and $Y$; the biased V-statistic is calibrated by a Gamma fit. HCR maps marginals to the copula domain via EDF and uses an orthonormal Legendre basis of degree $m$; we report (i) permutation $p$ for $T=\sum z^{2}$ ($\Sigma z^{2}$ perm”), (ii) permutation $p$ for $\max|z|$ (“max|z||z| perm”), and (iii) the $\chi^{2}$ approximation for $T$. Default settings: $m=6$, $B=800$ permutations.
(A) HSIC–Gamma $p$-values vs rotation; first rejection at $\approx 3.25^{\circ}$. (B) HCR $p$-values vs rotation for the three calibrations; max|z||z| perm rejects earliest ($\approx 2.75^{\circ}$), $\Sigma z^{2}$ perm and $\chi^{2}$ cross later ($\approx 3.75^{\circ}$). (C) Overlay of HSIC–Gamma and HCR on the same samples; all methods agree and decay with angle. (D) HSIC bandwidth sweep: RBF $\sigma$ set to median vs $2\timesmedian$; larger bandwidth is slightly more conservative, confirming the median heuristic as a reasonable default. (E) HSIC—zoom near $\alpha$ (detail of the crossing). (F) HCR—zoom near $\alpha$ for all three calibrations. (G) First-reject timeline ($p<\alpha$): horizontal markers summarize the angle of the first rejection for each method (earlier $\Rightarrow$ higher sensitivity). (H) HCR degree sweep (asymptotic $\chi^{2}$): comparison of $m\in\{4,6,8\}$ shows a modest gain from higher $m$ with stable behavior. Overall, increasing global rotation induces dependence and all tests detect it; HCR (max|z||z| perm) is the most sensitive here, followed by HSIC–Gamma, while $\Sigma z^{2}$ perm and $\chi^{2}$ calibrations are slightly more conservative but track the same trend.
IV Summary and further work
There was proposed application of HCR mutual information evaluation for independence test, and comparison with state-of-art HSIC. In much lower complexity, finally practical also to find subtle dependencies between very large data samples, it usually obtains even better sensitivity, additionally providing evaluation of mutual information, and interpretable description of the found dependencies, allowing to model joint distribution between data samples. We can further predict conditional probabilities - e.g. using HCR techniques from [14, 15].
As further work, the proposed tests likely can be improved, also from theoretical perspectives, choice of details especially of the final normality test - improving sensitivity and reducing false positive probability, e.g. Fig. 5 uses some standard ones.
There is also mentioned promising research direction of basis optimization e.g. to make it more likely for pairwise test to find dependencies - like initial PCA of data samples, or SVD of $a_{\textbf{j},\textbf{k}}$ matrix to optimize basis. However, it also amplifies noise - should be used on a family of similar samples to optimize basis for the given task, or for single sample using modified independence test e.g. based on random matrix theory.
We should also search for applications, like finding and exploiting subtle dependencies, or information bottleneck neural network training. Also compare with state-of-art mutual information estimation like [9].
Author contributions: Jagoda Bracha and Adrian Przybysz are responsible for Section III about HSIC and its tests with library, Jarek Duda is responsible for the rest.
Acknowledgments: The presented research was carried out within the AIntern program (https://www.qaif.org/events/aintern) organized by Fundacja Quantum AI, for project ”Biology-inspired artificial neurons with a joint distribution model”.

### v2

@@
 However, for size $n$ data sample it requires multiplication of $n\times n$ matrices, what currently needs $\sim O(n^{2.37})$ computational complexity [2], making it impractical for large data samples.
-We discuss HCR (Hierarchical Correlation Reconstruction) as its linear cost practical alternative of even higher dependence sensitivity in tests, and additionally providing actual joint distribution model by description of dependencies through features being mixed moments, starting with correlation and homoscedasticity, also allowing to approximate mutual information as just sum of squares of such nontrivial mixed moments between two data samples.
+We discuss HCR (Hierarchical Correlation Reconstruction) as its linear cost practical alternative, in tests of even higher sensitivity to dependencies, and additionally providing actual joint distribution model for chosen significance level, by description of dependencies through features being mixed moments, starting with correlation and homoscedasticity.
+Also allowing to approximate mutual information as just sum of squares of such nontrivial mixed moments between two data samples.
 Such single dependence describing feature is calculated in $O(n)$ linear time.
@@
 Statistical dependencies start with correlation coefficients, defining dependence between expected values of two random variable: their first moments.
-There are also higher moments, like variance, skewness, kurtosis - which also contribute to statistical dependencies, like homoscedasticity evaluating co-occurrence of high variance e.g.
-in financial time series.
-Figure 1: Proposed HCR independence test and comparison with state-of-art HSIC having much higher complexity, impractical for large samples.
+There are also higher moments, like variance, skewness, kurtosis - which also contribute to statistical dependencies, like homoscedasticity evaluating co-occurrence of high variance common e.g.
+in financial time series or Fig.
+1.
+Figure 1: Proposed HCR independence test and comparison with state-of-art HSIC of much higher complexity, impractical for large samples, which are required to find very subtle dependencies.
 We generate shown $n=1000$ size random dataset from Gaussian mixtures independently for both coordinates, and slightly rotate it to introduce dependence.
-In HCR we first separately normalize coordinates to nearly uniform distribution in $[0,1]$ with CDF/EDF, then model joint density as a chosen degree polynomial, here $m=4$.
+In HCR we first separately normalize coordinates to nearly uniform distribution in $[0,1]$ by CDF/EDF, then model joint density as a chosen degree polynomial, here $m=4$.
 Using orthonormal basis of Legendre polynomials, coefficients are approximately: expected value, variance, skewness, kurtosis, and their dependencies.
-Sum of squares of shown $m^{2}$ nontrivial coefficients approximate mutual information, and allows for independence test: by first normalizing multiplying by $\sqrt{n}$, and then testing if they agree with $N(0,1)$ set of variables, where the presented test of extreme coefficients has turned out the most sensitive to dependencies.
+Sum of squares of shown $m^{2}$ nontrivial coefficients approximate mutual information, and allows independence test: by first normalizing multiplying by $\sqrt{n}$, and then testing if they agree with $N(0,1)$ set of variables, where the presented test of extreme coefficients has turned out the most sensitive to dependencies, also evaluating significance of contributions to joint distribution.
 There is shown probability of exceeding these values with min/max from $m^{2}=16$ independent $N(0,1)$ variables, also both by taking their product using shown Mathematica code, properly indicating dependence with much higher certainty than HSIC, additionally providing its description with mixed moments and model of joint density.
 Hierarchical Correlation Reconstruction (HCR, [3, 4]) automatically decomposes dependencies into such mixed moments of two or more variables.
-Specifically, like in copula theory [5], we start with normalization of all variables to nearly uniform distribution in $[0,1]$ ($\sim$ quantiles) by $\hat{x}=\textrm{CDF}(x)$ or EDF (cumulative/empirical distribution function) separately for each coordinate.
+Specifically, like in copula theory [5], we start with normalization of all variables to nearly uniform distribution in $[0,1]$ ($\sim$ quantiles) by $\hat{x}=\textrm{CDF}(x)$ or EDF (cumulative/empirical distribution function), separately for each coordinate.
 Then we model joint density of such normalized variables as a linear combination in product basis:
@@
 For any continuous joint distribution we can approximate it with polynomials as close as we want, the $a_{\textbf{j}}$ coefficients become moments.
-Using orthonormal basis of polynomials $\int_{0}^{1}f_{k}(x)f_{l}(x)dx=\delta_{kl}$, coefficients can be MSE estimated as $a_{\textbf{j}}=\frac{1}{n}\sum_{\textbf{x}}f_{\textbf{j}}(\textbf{x})$ mean over normalized sample, in $O(n)$ time for $|X|=n$ data sample.
+Using orthonormal basis of (Legendre) polynomials $\int_{0}^{1}f_{k}(x)f_{l}(x)dx=\delta_{kl}$, coefficients can be MSE estimated as $a_{\textbf{j}}=\frac{1}{n}\sum_{\textbf{x}}f_{\textbf{j}}(\textbf{x})$ mean over normalized sample, in $O(n)$ time for $|X|=n$ data sample.
 Thanks to orthonormality, we can approximate mutual information as sum of squares of such nontrivial moments between two variables [4]: $I(X,Y)\approx\sum_{\textbf{j},\textbf{k}\neq\textbf{0}}(a_{\textbf{j},\textbf{k}})^{2}$.
 As independent variables have zero mutual information, in this article we will adapt it for independence test, asking for probability if all tested nontrivial moments between two variables could be zero.
-Alternative way to test independence is verifying if we can factorize joint distribution: $P_{X,Y}=P_{X}P_{Y}$?.
+Alternative way to test independence is verifying if we can factorize joint distribution: $P_{X,Y}=P_{X}P_{Y}$?
 It is used e.g.
@@
 Therefore, we discuss replacement of $\sim O(n^{2.37})$ HSIC independence test, with practical also for large samples: $O(n|B|)$ cost test based on HCR, for $B$ being a chosen set of features as mixed moments to test, e.g.
-growing $|B|\sim O(d^{2})$ with square of dimension if including only pairwise dependencies, or $O(d^{2})$ if adding triple-wise and so on.
+growing $|B|\sim O(d^{2})$ with square of dimension if including only pairwise dependencies, or $O(d^{3})$ if adding very subtle triplewise and so on.
+Beside much lower cost, making it practical to search very large data samples required to find subtle dependencies, HCR also turns out more sensitive to dependencies.
+It additionally provides approximation of mutual information, and model of joint distribution with control of significance of contributions.
 The original motivation of this article ([4]) was application for information bottleneck training of neural networks [6], planned as future work.
@@
 2, hopefully leading to improvements of information bottleneck training.
-This is initial version of article, with main purpose to introduce the method.
-We plan to extend benchmarks comparing with HSIC library [^1] HSIC library used for benchmarks: http://pypi.org/project/PyRKHSstats/ and details in the next version.
-Figure 2: Comparison of joint density estimation using global basis of polynomials/moments we use in HCR and local basis of Gaussians usually used in HSIC, requiring to choose radius $\epsilon$.
+This is early version of article, with main purpose to introduce the method.
+We plan to extend benchmarks comparing with HSIC library [^1] HSIC library used for benchmarks: http://pypi.org/project/PyRKHSstats/ and details in future versions.
+Figure 2: Comparison of joint density estimation using global basis of polynomials/moments we use in HCR, and local basis of Gaussians usually used in HSIC, requiring to choose radius $\epsilon$.
 Global basis allows to represent the density with a reasonable number of features (moments) e.g.
-$m^{2}$ here, allowing the discussed complexity reduction.
-Evaluating log-likelihood in cross-validation, local basis barely got above 0 obtained also for trivial $\rho=1$ assumption, while global basis was able to find features as moments, which generalize well from training to test set.
-Here with the highest log-likelihood for $m=4$ degree up to kurtosis, we focus on in this article.
-Intuitively, global features like moments are better at generalization, than local features just assuming that new points will be close to the old points.
+$m^{2}$ here, allowing the discussed reduction of complexity e.g.
+from $\sim O(n^{2.37})$ to $O(m^{2}n)$ for $n$ point independence test.
+Evaluating log-likelihood in cross-validation, local basis barely got above 0 obtained for trivial $\rho=1$ assumption, while global basis finds features as moments, which generalize well from training to test set.
+Here with the highest log-likelihood for $m=4$ degree up to kurtosis, quite universal e.g.
+for financial data, we focus on in this article.
+In contrast, local basis just assumes new points will be close to the old ones, what does not generalize well.
+Figure 3: Top left: the first $j=0,1,2,3,4$ from orthonormal polynomial basis $f_{j}(x)$ for $d=1$ dimension and product basis $f_{j}(x)f_{k}(y)$ for $d=2$ dimensions, which linear combination is used as model of (joint) density in HCR, e.g.
+$\rho(x,y)=\sum_{jk=1}^{m}a_{jk}f_{j}(x)f_{k}(y)$ for variables normalized to nearly uniform distribution in $[0,1]$.
+As $f_{0}=1$, top row $a_{i0}$ describes marginal distribution of $X$ and $a_{0j}$ left column of $Y$.
+Then $i,j\geq 1$ describe their dependencies as mixed moments close to expected value, variance, skewness and kurtosis.
+Right: $d=3$ case adding much more subtle triplewise dependencies.
+Bottom left: standard view on (joint/conditional) entropy of $X,Y$ variables and their mutual information.
+Right: for HCR using $\ln(1+t)\approx t$ first Taylor term they can be approximated with sums of squares of coefficients, especially mutual information as $I(X;Y)\approx\sum_{i,j=1}^{m}\left(a_{ij}\right)^{2}$, also in higher dimensions, which should be close to zero for independent variables.
 II HCR independence test and evaluation
-Assume we have $Z=\{(\textbf{x}^{i},\textbf{y}^{i})\}_{i=1}^{n}$ as size $n$ data samples as vectors of $d_{x}$, $d_{y}$ dimensions, for which we would like to test independence.
+Assume we have $X=\{\textbf{x}^{i}\}_{i=1}^{n},Y=\{\textbf{y}^{i}\}_{i=1}^{n}$ size $n$ data sample as pairs forming $Z=\{(\textbf{x}^{i},\textbf{y}^{i})\}_{i=1}^{n}$ of vectors of $d_{x}$, $d_{y}$ dimensions, for which we would like to test independence.
 To simplify presentation we focus on $d_{x}=d_{y}=d$ equal dimension case, but also discuss general situation.
-In HCR it is convenient to start with normalization of marginal distributions to nearly uniform in $[0,1]$, then model their joint distribution as a linear combination in some orthonormal basis, usually polynomials - which should use zero coefficients for independent variables, we will build independence test from.
+In HCR it is convenient to start with normalization of marginal distributions to nearly uniform in $[0,1]$, then model joint distribution as a linear combination in some orthonormal basis, usually polynomials - which for independent samples should use zero nontrivial moments between them, we can build test from.
 II-A Normalization to nearly uniform in $[0,1]$ with CDF/EDF
-As in copula theory citecopula, we start with normalization of variables to nearly uniform distribution in $[0,1]$ by transforming with CDF/EDF (cumulative or empirical distribution function) to nearly quantiles, independently for each coordinate.
+As in copula theory  [5], we start with normalization of variables to nearly uniform distribution in $[0,1]$ by transforming with CDF/EDF (cumulative or empirical distribution function) to nearly quantiles, independently for each coordinate.
 For CDF approach we need to choose some parametric family e.g.
@@
 However, it requires sorting having $O(n\ln n)$ complexity - if linear is required, there should be used CDF.
-Figure 3: Top left: the first $j=0,1,2,3,4$ from orthonormal polynomial basis $f_{j}(x)$ for $d=1$ dimension and product basis $f_{j}(x)f_{k}(y)$ for $d=2$ dimensions, which linear combination is used as model of (joint) density in HCR, e.g.
-$\rho(x,y)=\sum_{jk=1}^{m}a_{jk}f_{j}(x)f_{k}(y)$ for variables normalized to nearly uniform distribution in $[0,1]$.
-As $f_{0}=1$, top row $a_{i0}$ describes marginal distribution of $x$ and $a_{0j}$ left column of $y$.
-Then $i,j\geq 1$ describe their dependencies as mixed moments close to expected value, variance, skewness and kurtosis.
-Right: $d=3$ case adding triplewise dependencies.
-Bottom left: standard view on (joint/conditional) entropy of $X,Y$ variables and their mutual information.
-Right: for HCR using $\ln(1+w)\approx w$ first Taylor term they can be approximated with sums of squares of coefficients, especially mutual information as $I(X;Y)\approx\sum_{i,j=1}^{m}\left(a_{ij}\right)^{2}$, also in higher dimension, which should be close to zero for independent variables.
 For the discussed independence test/evaluation in benchmarks we use EDF: for $d$-dimensional data samples $X$ and $Y$, we start with transforming them by applying EDF normalization separately for each coordinate as $\textrm{EDF}^{x}_{i},\textrm{EDF}^{y}_{i}$:
 $\hat{Z}=\{(\hat{\textbf{x}},\hat{\textbf{y}}):\hat{x}_{i}=\textrm{EDF}^{x}_{i}(x_{i}),\hat{y}_{i}=\textrm{EDF}^{y}_{i}(y_{i}),(\textbf{x},\textbf{y})\in Z\}$ (2)
-As mutual information is invariant under reparametrization of the marginal variables [9], we can estimate it for such transformed variables instead: $I(U;V)\approx I(X;Y)$.
+[…]
 II-B Polynomial density model
@@
 $f_{4}(x)=3\left(70x^{4}-140x^{3}+90x^{2}-20x+1\right)\qquad\sim\textrm{kurtosis}$
+The coefficients can be interpreted as mixed moments, only between nonzero indexes as $f_{0}=1$, e.g.
+$a_{00i0}$ is $i$-th moment of 3rd out of 4 variables, $a_{012}$ describes dependence between expected value of 2nd variable, and variance of 3rd variable.
 Alternatively there could be used Fourier or DCT basis, especially for periodic variables.
-The coefficients can be interpreted as mixed moments, as $f_{0}=1$ only between nonzero indexes, e.g.
-$a_{00i0}$ is $i$-th moment of 3rd out of 4 variables, $a_{012}$ describe dependence between expected value of 2nd variable, and variance of 3rd variable.
+As mentioned further, the basis can be automatically optimized e.g.
+with SVD [10], for given data sample, or family of data.
 II-C Entropy, mutual information estimation
-Using $\ln(1+t)\approx t$ approximation and orthonormality, we can approximate entropy in nits ($1/\ln(2)\approx 1.44$ bits) from the coefficients:
+Using $\ln(1+t)\approx t$ first order approximation and orthonormality, we can approximate entropy in nits ($1/\ln(2)\approx 1.44$ bits) from the coefficients:
 $H(\hat{X})=-\int_{[0,1]^{d}}\rho(\mathbf{x})\,\ln(\rho(\mathbf{x}))\,d\mathbf{u}\approx-\sum_{\mathbf{j}\in B_{x}^{+}}(a_{\mathbf{j}})^{2}$ (6)
@@
 However, using these sum of squares formulas with (4) estimator as mean: $a_{\textbf{j},\textbf{k}}=\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))$, it would be artificially increased - let us now discuss it and correction.
-II-C1 Subtracting the variance
+Subtracting the variance
 Assume we get $n$ values $\{w^{i}\}_{i=1..n}$ from $N(\mu,\sigma)$ distribution, and we want to estimate $\mu^{2}$.
-A natural approach is calculating the mean $\bar{w}=\frac{1}{n}\sum_{i}w_{i}$ and taking $\bar{w}^{2}$, however, it would be artificially increased.
+A natural approach is calculating the mean $\bar{w}=\frac{1}{n}\sum_{i}w_{i}$ and taking $\bar{w}^{2}$, however, it would be biased (artificially increased).
 The Central Limit Theorem (CLT) says that for large sample:
 $\textrm{CLT}:\qquad\qquad\frac{\bar{w}-\mu}{\sigma/\sqrt{n}}\sim N(0,1)$ (9)
-Applying square to both sides, and taking expected values: mean square from $N(0,1)$ is 1, hence we get:
-$n\,E[(\bar{w}-\mu)^{2}/\sigma^{2}]=1$
+Applying square to both sides, and taking expected values: mean square of value from $N(0,1)$ is 1, hence we get:
+$\frac{n}{\sigma^{2}}\,E[(\bar{w}-\mu)^{2}]=1$
 As $E[\bar{w}]=\mu$, expanding the square we finally get
@@
 […]
-$I_{c}(\hat{X},\hat{Y})=\sum_{\mathbf{j}\in B_{x}^{+}}\ \sum_{\mathbf{k}\in B_{y}^{+}}\left(\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))^{2}-\frac{1}{n}\textrm{var}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))\right)$ (11)
+$I_{c}(\hat{X},\hat{Y})=\sum_{\mathbf{j}\in B_{x}^{+}}\ \sum_{\mathbf{k}\in B_{y}^{+}}\left(\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))^{2}-\frac{\textrm{var}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))}{n}\right)$ (11)
 we plan to compare in future with state-of-art methods like [9].
+Also, while we have used only the first term of Taylor series $\ln(1+t)=-\sum_{i=1}^{\infty}(-t)^{i}/i$ , it is worth to somehow include higher terms, requiring integrals of 3 and more basis functions.
 II-D HCR-based independence tests
 […]
-To test it, let us first normalize them as in CLT:
+To test it, let us first normalize them as in CLT, so that for $H_{0}$ independence hypothesis, all $\hat{a}_{\textbf{j},\textbf{k}}$ should be from $N(0,1)$:
 $M=|B^{+}_{x}||B^{+}_{y}|\qquad\textrm{of}\qquad\hat{a}_{\textbf{j},\textbf{k}}=\frac{\textrm{mean}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))}{\sqrt{\textrm{var}(f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{y}))/n}}$ (12)
-Assuming $H_{0}$ hypothesis, for pairwise dependencies: between one coordinate of both samples, due to independence and normalization above ’var’ variance has to be 1, allowing to omit the above estimation from sample.
+Assuming $H_{0}$ hypothesis, for pairwise dependencies: between one coordinate of each samples, due to independence and normalization above ’var’ variance has to be 1, allowing to omit the above estimation from sample.
 However, for higher order dependencies it does not have to be true due to dependencies inside samples - requiring to estimate variance from sample.
-Accordingly to CLT, for $H_{0}$ independence hypothesis, all $\hat{a}_{\textbf{j},\textbf{k}}$ should be from approximately $N(0,1)$.
 […]
-II-D1 Log-likelihood test
-of $\textrm{mean}(\ln(\textrm{PDF}(\hat{a}_{\textbf{j},\textbf{k}})))$, which assuming $H_{0}$ and large $M$ is approximately from normal distribution.
-For $N(0,1)$ PDF we get $E[\ln(\textrm{PDF})]=-(1+\ln(2))/2\approx-1.4189$ and variance 1/2, leading to test:
+Log-likelihood test
+of $\textrm{mean}(\ln(\textrm{PDF}(\hat{a}_{\textbf{j},\textbf{k}})))$, which assuming $H_{0}$ and large $M$ is approximately from $N(0,1)$, for which we get $E[\ln(\textrm{PDF})]=-(1+\ln(2))/2\approx-1.4189$ and variance 1/2, leading to test:
 $\textrm{Pr}(H_{0})=1-\textrm{erf}\left(\frac{|\textrm{mean}(\ln(\textrm{PDF}(\hat{a}_{\textbf{j},\textbf{k}})))+1.4189|}{2\sqrt{M}}\right)$ (13)
-Figure 4: $\textrm{PDF}^{M}_{i}$ probability densities of sorted $M=1,\ldots,16$ variables from $N(0,1)$ normalized Gaussian distribution, which can be used for tests if estimated normalized coefficients are from $N(0,1)$ for $\mathcal{H}_{0}$ hypothesis of independence, especially for the extreme values indicating the main dependence direction.
-II-D2 Testing extremes $\{\hat{a}_{\textbf{j},\textbf{k}}\}_{\textbf{j}\in B_{x}^{+},\textbf{k}\in B_{y}^{+}}$
+Figure 4: $\textrm{PDF}^{M}_{i}$ probability densities of sorted $M=1,\ldots,16$ variables from $N(0,1)$ normalized Gaussian distribution (analogously for different distributions), which can be used for tests if estimated normalized coefficients are from $N(0,1)$ for $\mathcal{H}_{0}$ hypothesis of independence, especially for the extreme values indicating the main dependence direction.
+As in Fig.
+5, it also allows to estimate significance of individual contributions, e.g.
+for joint distribution model of chosen significance level.
+Testing extremes $\{\hat{a}_{\textbf{j},\textbf{k}}\}_{\textbf{j}\in B_{x}^{+},\textbf{k}\in B_{y}^{+}}$
 […]
-For hypothesis testing, we need to find distribution for $M$ sorted values from $N(0,1)$.
+For hypothesis testing, we need to find distribution of sorted $M$ values from independent $N(0,1)$ or a different single variable distribution having given $\textrm{PDF}=\textrm{CDF}^{\prime}$, e.g.
+for Marchenko-Pastur for SVD optimization.
 To find $\textrm{PDF}_{M,i}(x)$ density of $i$-th in order, there are i1i-1 smaller or equal, and M1iM-1-i greater then equal, combinatorially leading to $M{M-1\choose i-1}$ possibilities, with density, shown in Fig.
@@
 $\textrm{CDF}^{M}_{i}(x)=M{M-1\choose i-1}\,\textrm{Beta}(\textrm{CDF},i,1-i+M)$ (15)
-using incomplete Euler Beta function. We can use $\textrm{CDF}_{M,i}$ to test probability of obtaining one or a few extreme values.
-II-E Choice of basis $B$
+using incomplete Euler Beta function.
+We can use $\textrm{CDF}^{M}_{i}$ to test sorted values for hypothesis of being from a given distribution, also find those disagreeing for a given significance level.
+II-E Chosen statistical significance joint distribution model
+[…]
+$\min\left(\textrm{CDF}^{M}_{i}(\textrm{sort}(\hat{a})_{i}),\ 1-\textrm{CDF}^{M}_{i}(\textrm{sort}(\hat{a})_{i})\right)<\alpha$ (16)
+for $N(0,1)$ distribution in joint density formula (3), or Marchenko-Pastur-like for SVD optimized as in Fig. 5.
+Figure 5: Top: we can perform SVD (singular value decomposition) of $|B_{x}^{+}|\times|B_{y}^{+}|$ matrix of coefficients, allowing to optimize their bases (18), leading to decomposition of modeled joint density into contributions of controlled significance, allowing to use the highest or a few below chosen significance level $\alpha$, here leading to similar Pr(independent) as HCRboth in Fig.
+1.
+To distinguish signal from noise, we could use random matrix theory for singular values of $|B_{x}^{+}|\times|B_{y}^{+}|$ matrix with $N(0,1)$ random coefficients.
+Marchenko-Pastur theorem [13] handles such situations, however, for the limit of infinite size, e.g.
+bounding singular values by 4.
+In contrast, the plots shows densities from $10^{6}$ generated $4\times 4$ random matrices of $N(0,1)$ coefficients, clearly exceeding this bound - we can use such empirical estimation for practical approximation.
+It allows for joint distribution models using statistically significant contributions, e.g.
+for 0.01 significance we should take normalization plus the first two.
+Bottom: analogously directly using coefficients for various significance level using (16) formula.
+There are also shown $\textrm{PDF}^{M}_{i}$ densities for $M=16$ sorted $N(0,1)$ values, and sorted $\hat{a}$ coefficients being $N(0,1)$ for independence hypothesis, showing that non-extreme values can also be significant.
+II-F Choice of basis $B$
 For $d_{x}=d_{x}=d=1$ dimensional samples, a natural choice of basis is just $B^{+}=\{1,2,\ldots,m\}$ for some $m$ number of moments, e.g.
@@
 2, can be optimized for some specific data this way, or e.g.
-trying to recognize small rotations for like in Fig.
+trying to recognize small rotations like in Fig.
 1.
@@
 Dependence between samples should be usually detectable through such pairwise dependencies.
-However, in theory there can be only more subtle higher order dependencies - HCR independence test would need to include to be able to detect, evaluate, describe.
+However, in theory there can be only more subtle higher order dependencies - HCR independence test would need to consciously include to be able to detect, evaluate, describe.
 For example by adding to $B_{x}$ and/or $B_{y}$ basis $\sum_{i}\textrm{sign}(j_{i})=2$ indexes, e.g.
 $a_{11,1}$ would include triplewise dependence like in Fig.
-3: that with change of both coordinates, there is change of expected value in the third one.
-The cost is larger number of features: $O(d^{2})$ for pairwise, $O(d^{3})$ for triplewise, and so on.
+3: that with change of two coordinates, there is change of expected value of the third one.
+The cost is larger number of features to estimate and analyze: $O(d^{2})$ for pairwise, $O(d^{3})$ for triplewise, and so on.
+II-G PCA, CCA, SVD optimizations
 We can also perform some preprocessing of data to try to emphasize dependencies, hopefully making it more likely for low order e.g.
 pairwise test to find them, like PCA (principal component analysis) - rotation of coordinates to eigenvectors of correlation matrix, separately for both samples.
-Also e.g.
-using SVD (singular value decomposition) of $a_{\textbf{k},\textbf{j}}$ as $|B^{+}_{x}|\times|B_{y}^{+}|$ matrix, we can optimize both used bases.
-We can freely perform such optimization working on a family of data samples of similar behavior.
-However, performing such optimization separately on a single tested sample would also amplify its noise - it would need modification of independence test to include it, e.g.
+PCA optimization often amplifies noise, e.g.
+Canonical Correlation analysis (CCA) allows to repair it - optimizing basis to maximize correlations, as discussed for HCR in [14].
+For independence test we could use it to maximize correlations between $f_{\textbf{j}}(\textbf{x})$ and $f_{\textbf{k}}(\textbf{y})$ values in bases for both samples.
+For final $a_{\textbf{j},\textbf{k}}$ as $|B^{+}_{x}|\times|B_{y}^{+}|$ matrix we can also use SVD (singular value decomposition) into $U\Sigma V^{T}$ for orthogonal $U^{T}U=V^{T}V=I$ and diagonal $\Sigma$ of singular values, like in Fig.
+5 we can optimize both used bases $f_{B_{x}}(\textbf{x})=(f_{\textbf{j}}(\textbf{x}))_{\textbf{j}\in B_{x}}$, $f_{B_{y}}(\textbf{x})=(f_{\textbf{k}}(\textbf{y}))_{\textbf{k}\in B_{y}}$ multiplying them by $U$ and $V$, to represent joint density:
+$\rho(\textbf{x},\textbf{y})=(f_{B_{x}}(\textbf{x}))(a_{\textbf{j},\textbf{k}})(f_{B_{y}}(\textbf{y}))^{T}=\left((f_{B_{x}}(\textbf{x}))U\right)\Sigma\left((f_{B_{y}}(\textbf{y}))V\right)^{T}$ (18)
+We can freely perform such optimizations working on a family of data samples of similar behavior, using PCA, CCA or SVD of e.g.
+averaged samples, matrices.
+However, performing such optimization separately on a single tested sample would also amplify its noise - like in Fig.
+5, it needs modification of independence test to include it, e.g.
 based on random matrix theory, or synthetic empirical simulations.
-II-F HCR independence test used in benchmarks
+II-H HCR independence test used in benchmarks
 The benchmarks for HCR have used basic pairwise dependencies for fixed degree $m=4$, testing minimal and maximal considered $\hat{a}$, presented as Algorithm 1 and in Fig. 1.
@@
 […]
-Accepting $H_{0}$ hypothesis if returned Pr(independent) is above some chosen significance level, or reject otherwise.
+Accepting $H_{0}$ hypothesis if returned Pr(independent) is above some chosen significance level $\alpha$, or reject otherwise.
 It uses $\textrm{var}=1$ in (12) as mentioned allowed for pairwise dependencies, for triplewise and higher $\hat{a}$ should be additionally divided by estimated variance for coefficients $\textbf{j},\textbf{k}$ having two or more nonzero indexes.
@@
 It is also worth to consider different final tests if all $M$ normalized features are from $N(0,1)$ distribution - testing min and max is very sensitive to dependencies, but could lead to false positives, and ignores intermediate values.
-It can be extended from $2$ to $2k$ extreme values by sorting $\hat{a}$ to $\textrm{Pr}(\textrm{independent})\approx$
+As we can see in Fig.
+5, also non-extreme values can be valuable to distinguish signal from noise, suggesting to extend independence test to even more sensitive: by using $2k\leq M$ extreme values, $i$-th sorted $\hat{a}$ as $\textrm{sort}(\hat{a})_{i}$ for $\textrm{Pr}(\textrm{independent})\approx$
 $2^{2k}\prod_{i=1}^{k}\textrm{CDF}_{i}^{M}\left(\textrm{sort}(\hat{a})_{i}\right)\,\left(1-\textrm{CDF}_{M+1-i}^{M}\left(\textrm{sort}(\hat{a})_{M+1-i}\right)\right)$ (19)
-Where $2^{2k}$ is rescaling to get approximately 1 for independent. To include all values we could use e.g. Anderson-Darling, Shapiro-Wilk tests, $\chi^{2}$, permutation, or log-likelihood test.
-II-G Toward kernel methods: global HCR vs local HSIC basis
+Where $2^{2k}$ is rescaling to get approximately 1 for independent.
+Analogously, as in Fig.
+5, we can perform such independence test after SVD basis optimization, comparing singular values of $\hat{a}_{\textbf{j},\textbf{k}}$ instead of the original coefficients.
+To include all the values we could also use e.g. Anderson-Darling, Shapiro-Wilk, $\chi^{2}$, permutation, or log-likelihood test.
+II-I Toward kernel methods: global HCR vs local HSIC basis
 As in [4], we can represent the features as $n\times|B|$ matrices:
@@
   • global basis is usually better at generalization: provides higher log-likelihood in cross-validation, as describing general features, instead of just assuming that new points will be close to the old points for local basis,
-  • global basis often allows to describe density with a reasonable number of features (moments), allowing more practical description e.g. fo,x’)r the discussed reduced complexity,
+  • global basis often allows to describe density with a reasonable number of features (moments), allowing more practical description e.g. for the discussed reduced complexity,
   • local basis requires to choose kernel width, while global to choose basis with often universal e.g. $m=4$ degree,
   • global orthonormal basis for normalized variables can work on deformation of $\rho=1$, allowing for practical approximations of entropy, mutual information using $\ln(1+t)\approx t$.
-  • above HCR kernel matrices include all $B_{x}\times B_{y}$ features, while for HSIC Gaussian kernel only single $\sigma$ is used:
+  • above HCR kernel matrices include all $|B_{x}^{+}||B_{y}^{+}|$ features, while for HSIC Gaussian kernel only single $\sigma$ is used:
 $K^{\textrm{HCR}}_{\textbf{x},\textbf{x}^{\prime}}=\frac{1}{n}\sum_{\mathbf{j}\in B_{x}^{+},\mathbf{k}\in B_{x}^{+}}f_{\textbf{j}}(\textbf{x})f_{\textbf{k}}(\textbf{x}^{\prime})\quad\qquad K^{\textrm{HSIC}}_{\textbf{x},\textbf{x}^{\prime}}=e^{-\frac{\|\textbf{x}-\textbf{x}^{\prime}\|_{2}}{2\sigma^{2}}}$
+maybe worth to extend e.g. to multiple $\sigma_{i}$: $\sum_{i}\exp\left(-\frac{\|\textbf{x}-\textbf{x}^{\prime}\|_{2}}{2\sigma_{i}^{2}}\right)$.
 III Hilbert–Schmidt Independence Criterion (HSIC)
@@
 In practice, they are complementary: HSIC for broadly powered detection, HCR for scalable detection plus explanation of the dependence structure.
+IV Summary and further work
+There was proposed application of HCR mutual information evaluation for independence test, and comparison with state-of-art HSIC.
+In much lower complexity, finally practical to also work with very large data samples - crucial to find very subtle dependencies, it usually obtains even better sensitivity.
+Additionally, HCR provides evaluation of mutual information, and interpretable description of the found dependencies, allowing to model joint distribution between data samples, including significance evaluation of contributions.
+Finding dependencies, We can further exploit them to predict conditional densities - e.g.
+using HCR techniques from [17, 18, 14].
+As further work, the proposed tests likely can be improved, also from theoretical perspectives, choice of details especially of the final normality test - improving sensitivity and reducing false positive probability, e.g.
+Fig.
+6 uses some standard ones.
+The proposed test (19) of sorting and individually comparing multiple values with distributions of sorted (15) e.g.
+$N(0,1)$ seems promising general approach for normality test, also for other distributions - worth to be explored.
+There is also mentioned promising research direction of basis optimization e.g.
+to make it more likely for pairwise test to find dependencies - like initial PCA or CCA of data samples, or directly SVD of $a_{\textbf{j},\textbf{k}}$ matrix to optimize basis as in Fig.
+5.
+It also amplifies noise - should be used on a family of similar samples to optimize basis for the given task, or for single sample using modified independence test e.g.
+based on random matrix theory or simulations randomly generating matrices.
+Finally we should explore applications, like finding and exploiting subtle dependencies e.g.
+in financial data, or information bottleneck neural network training.
+Also compare with state-of-art mutual information estimation like [9].
+Author contributions: Jagoda Bracha and Adrian Przybysz are responsible for Section III about HSIC and its tests with library, Jarek Duda is responsible for the rest.
+Acknowledgments: The presented research was carried out within the AIntern program (https://www.qaif.org/events/aintern) organized by Fundacja Quantum AI, for project ”Biology-inspired artificial neurons with a joint distribution model”.
 Figure 6: HSIC (Gamma) vs HCR.
@@
 Overall, increasing global rotation induces dependence and all tests detect it; HCR (max|z||z| perm) is the most sensitive here, followed by HSIC–Gamma, while $\Sigma z^{2}$ perm and $\chi^{2}$ calibrations are slightly more conservative but track the same trend.
-IV Summary and further work
-There was proposed application of HCR mutual information evaluation for independence test, and comparison with state-of-art HSIC.
-In much lower complexity, finally practical also to find subtle dependencies between very large data samples, it usually obtains even better sensitivity, additionally providing evaluation of mutual information, and interpretable description of the found dependencies, allowing to model joint distribution between data samples.
-We can further predict conditional probabilities - e.g.
-using HCR techniques from [14, 15].
-As further work, the proposed tests likely can be improved, also from theoretical perspectives, choice of details especially of the final normality test - improving sensitivity and reducing false positive probability, e.g.
-Fig.
-5 uses some standard ones.
-There is also mentioned promising research direction of basis optimization e.g.
-to make it more likely for pairwise test to find dependencies - like initial PCA of data samples, or SVD of $a_{\textbf{j},\textbf{k}}$ matrix to optimize basis.
-However, it also amplifies noise - should be used on a family of similar samples to optimize basis for the given task, or for single sample using modified independence test e.g.
-based on random matrix theory.
-We should also search for applications, like finding and exploiting subtle dependencies, or information bottleneck neural network training.
-Also compare with state-of-art mutual information estimation like [9].
-Author contributions: Jagoda Bracha and Adrian Przybysz are responsible for Section III about HSIC and its tests with library, Jarek Duda is responsible for the rest.
-Acknowledgments: The presented research was carried out within the AIntern program (https://www.qaif.org/events/aintern) organized by Fundacja Quantum AI, for project ”Biology-inspired artificial neurons with a joint distribution model”.